# Saliency Guided Image Warping for Unsupervised Domain Adaptation

### v2

@@ Abstract
Driving is challenging in conditions like night, rain, and snow. The lack of good labeled datasets has hampered progress in scene understanding under such conditions. Unsupervised domain adaptation (UDA) using large labeled clear-day datasets is a promising research direction in such cases. Current UDA methods, however, treat all image pixels uniformly, leading to over-reliance on the dominant scene backgrounds (e.g., roads, sky, sidewalks) that appear dramatically different across domains. As a result, they struggle to learn effective features of smaller and often sparse foreground objects (e.g., people, vehicles, signs).
In this work, we improve UDA training by using in-place image warping to focus on salient object regions. Our insight is that while backgrounds vary significantly across domains (e.g., snowy night vs. clear day), object appearances vary to a lesser extent. Therefore, we design instance-level saliency guidance to adaptively oversample object regions, which reduces adverse effects from background context and enhances backbone feature learning. We then unwarp the better learned features while adapting from source to target. Our approach improves adaptation across geographies, lighting, and weather conditions, and is agnostic to the task (segmentation, detection), domain adaptation algorithm, saliency guidance, and underlying model architecture. Result highlights include +6.1 mAP50 for BDD100K Clear $\rightarrow$ DENSE Foggy, +3.7 mAP50 for BDD100K Day $\rightarrow$ Night, +3.0 mAP50 for BDD100K Clear $\rightarrow$ Rainy, and +6.3 mIoU for Cityscapes $\rightarrow$ ACDC. Our method adds minimal training memory and incurs no additional inference latency. Please see Appendix for more results and analysis.
Figure 1: Saliency Guided Image Warping improves domain adaptive object detection across a extensive range of lighting and weather scenarios in various target datasets (or subsets) using supervision from the BDD100K source subset containing day images in clear weather.
@@ 1 Introduction
Figure 2: Object-Background Pixel Imbalance & Differences in Cross-Domain Object-Background Variations. Consider scenes from different domains: clear day, rainy, and night. We use red to highlight a foreground object and yellow to highlight some background areas. Note, (a) background pixels occupy more space than foreground object pixels, (b) background elements like sky and road exhibit higher cross-domain variations, while foreground objects like cars show lesser variations. From these observations, we believe focusing more on objects would mitigate over-reliance on the background and improve domain adaptation.
@@
Camera-captured images, meanwhile, exhibit an object-background pixel imbalance, where pixel distribution is skewed towards backgrounds taking up more space. In contrast, objects are smaller (e.g. traffic signs), more scattered, and thus occupy fewer pixels. The camera’s perspective further amplifies this effect, making distant objects appear smaller and allowing larger backgrounds to dominate the scene. However, contemporary perception models treat image pixels, whether objects or backgrounds, uniformly. This leads to an over-reliance on the dominant backgrounds, making it challenging to learn effective features for less prevalent objects amid these dominant backgrounds.
@@
This over-reliance on background pixels becomes even more problematic in Unsupervised Domain Adaptation (UDA) [13, 14, 17], which is crucial for driving in challenging conditions due to the scarcity of labeled data. However, UDA exhibits differences in cross-domain object-background variations, where background regions show greater variations than object regions across domains. Consequently, this over-reliance on background pixels makes adaptation difficult. Reducing dependence on these variable backgrounds would improve adaptation to new domains.
@@
One possible remedy is to train at various image scales using data augmentation (e.g., multi-scale training [33] and multi-scale inference [11, 32, 14]). However, these strategies introduce significant computational and storage overhead. Another approach, employing scale invariance priors (e.g., feature pyramids [19, 22, 21] or test-time scale optimization [8, 4]), only partially improve recognition of different sizes of objects and introduces additional inference latency. Yet another strategy is to oversample or ‘zoom’ into salient image regions [24, 35, 7, 9, 36]. However, these methods do not distinguish between object and background, are only applicable to supervised settings, and introduce additional inference latency due to test-time operations.
@@
To address these limitations, we develop a novel strategy that judiciously utilizes the available pixel space in an image to focus on salient regions for unsupervised domain adaptation. We adaptively oversample salient image regions corresponding to objects in-place, directly warping the original image without creating new images or additional crops. Then, we unwarp the better learned features while adapting models from a source to a target domain. This reduces background context dependence for detection and segmentation, significantly improving domain adaptability. During inference, we do not perform any warping as the learned features are already improved, thus adding no latency. We show significant improvements on a large set of real to real adaptation scenarios compared to prior work [13, 17, 14, 15] (Sections 5.15.2, Appendix A, Appendix B).
In summary, our contributions are:
@@
  • We are the first to demonstrate that saliency guided image warping enhances backbone features and improves model adaptability for unsupervised domain adaptation (Figure 1, Section 5.2 and Appendix E).
@@
  • Our approach is efficient, with train-time image warping and feature unwarping but no test-time warping, resulting in minimal training overhead and no additional inference latency (Appendix B).
@@ 2 Related Work
Figure 3: Saliency Guided Image Warping for Unsupervised Domain Adaptation. Unsupervised Domain Adaptation algorithms (e.g., [13, 17]) adapt from the source domain to the target domain in two stages: supervised pre-training on the source domain and unsupervised adaptation from source to target. Components of warping and unwarping are marked in cyan. For a source domain image, we generate saliency guidance based on source bounding boxes (Section 3.2) and unwarp the features before computing the supervised loss in the unwarped label space. While guidance can be computed via Static Prior [35, 36] or Geometric Prior [9], our Instance-Level Guidance, which focuses on oversampling object regions, shows better performance and incurs no test-time latency (Sections 5.1 and 5.2). Meanwhile, the unsupervised losses computed on the target image by the Domain Adaptation Algorithm (e.g.,  [13, 17, 14, 15, 18, 6, 49]) remains the same.
@@
Saliency Guided Image Warping. Non-uniform Spatial Transformations have been used to correct image distortions arising from camera acquisition [43]. For example, several spatial transformation methods [16, 24] for learned models introduce intentional image distortions to enhance performance. Learning to Zoom [24] showed that over-sampling task-specific salient regions is useful, as visual recognition requires combining information from different spatial resolutions. Building on this paradigm, warping for other tasks like detection [35, 36] and segmentation [36] has been explored. For instance, Fovea [35] and Two Plane Prior [9] designed ‘salient’ priors to guide warping, introducing static & temporal, and geometric saliency guidance, respectively.
@@
Our approach stands outs in the following ways: (a) Unlike former image warping methods that indiscriminately process object and background, we introduce an analytical instance-level saliency guidance that targets regions containing objects. (b) Unlike previous image warping efforts focused on supervised settings, our novel training strategy and saliency guidance tailor image warping for unsupervised domain adaptation. (c) Unlike prior warping works like [35, 9, 36], our strategy introduces no additional inference latency as we do not warp during testing.
@@
Image Transformations for Domain Adaptation. Image transformations like flipping, cropping, blurring, color jittering, grayscaling, and contrast adjustment are commonly used as perturbations to enhance domain adaptation [38, 46, 47, 23, 29, 49, 37, 6, 41, 44, 18, 13, 14, 12, 15, 17]. These techniques increase source domain diversity and narrow the gap to the target domain. For example, 2PCNet [17] introduces a domain-specific Night Augmentation (NightAug) to make day images resemble night images. To the best of our knowledge, in-place image warping as a domain adaptive perturbation has not been explored previously. Our approach is agnostic to domain adaptation algorithms and can be seamlessly combined with other augmentation strategies.
Figure 4: Image Warping with Different Saliency Guidance Functions. In-place warping follows a zero-sum pixel constraint: enlarging one region necessitates shrinking another. Bounding boxes mark small, medium, and large objects in the image. Our instance-level saliency guidance oversamples objects and undersamples the background. In contrast, the Static Prior fails when object locations do not align with the dataset’s average object location, while the Geometric Prior fails for small objects not near the vanishing point.
@@ 3 Method
We first describe how to warp images and unwarp features using a spatial guidance function. Next, we introduce our novel warping guidance to explicitly oversample objects (as shown in Figure 1). Finally, we discuss incorporating our warping strategy into self-training unsupervised domain adaptation algorithms (See Figure 3).
@@ 3.1 Image Warping and Feature Unwarping
Image Warping: We consider an inverse transformation $\mathcal{T}^{-1}_{S}$ parameterized by saliency $S$ such that:
@@
a resampled $I^{\prime}$ for every location $\mathbf{u}$ is obtained by using warp $W_{\mathcal{T}}$ given input image $I$ and saliency guidance $S$, which is a 2D map of the output image size. The rationale and choice of $S$ is described in Section 3.2.
Feature Unwarping: If we operate in the space of warped images, our predictions would also be warped. However, we do not wish to warp our labels to match the warped input image, Instead, we perform feature unwarping of spatial backbone features and train the model without warping our labels or unwarping our predictions like [35, 9]. This strategy makes our warping mechanism task-agnostic. We adopt the strategy from LZU [36], which approximates $\mathcal{T}$ with $\widetilde{\mathcal{T}}$, a piecewise tiling of invertible bilinear maps, and then perform feature unwarping via:
${\mathcal{T}}^{-1}(\mathbf{u})\approx\widetilde{\mathcal{T}}^{-1}(\mathbf{u})=\begin{cases}\widetilde{\mathcal{T}}_{ij}^{-1}(\mathbf{u})&\text{ if }\mathbf{u}\in\operatorname{Range}\left(\widetilde{\mathcal{T}}_{ij}\right)\\ 0&\text{ else }\end{cases}$ (2)
Additional Considerations. We restrict the space of allowable warps for $\mathcal{T}$ to preserve axis alignment. In this approach, entire rows or columns are ‘stretched’ or ‘compressed,’ which are shown by [35, 9, 36] to perform better:
$\mathcal{T}_{S,u}^{-1}(u)=\frac{\int_{u^{\prime}}S_{u}(u^{\prime})k(u^{\prime},u)u^{\prime}}{\int_{u^{\prime}}S_{u}(u^{\prime})k(u,u^{\prime})}$ (3)
$\mathcal{T}_{S,v}^{-1}(v)=\frac{\int_{v^{\prime}}S_{v}(v^{\prime})k(v^{\prime},v)v^{\prime}}{\int_{v^{\prime}}S_{v}(v^{\prime})k(v,v^{\prime})}$ (4)
where $k$ is a Gaussian kernel and $\mathbf{u}=(u,v)$ represent the two image axes in pixel space. $S_{u}$ and $S_{v}$ are the marginalized versions of $S$ along $u$ and $v$ axes, respectively.
@@ 3.2 Designing Saliency Guidance
Saliency ($S$) determines which regions of an image are sampled more during warping [24]. We explore three saliency functions: two based on approximate guidance from previous work [35, 36, 9], and a novel instance-level saliency guidance proposed in this work. Since images are warped in-place, expanding some regions necessitates reducing others (Figure 4). Only our instance-level saliency guidance oversamples objects and undersamples the background.
Figure 5: Domain Adaptive Object Detection (BDD100K Day $\rightarrow$ Night). Model trained with our method demonstrate an improved scene understanding ability to distinguish between different objects and between objects and backgrounds. In contrast, 2PCNet [17] misclassifies roads as cars and streetlights as traffic signs.
Approximate Saliency Guidance: The Geometric Prior [9] saliency guidance is a function of the dominant vanishing point in the scene, oversampling objects close to the vanishing point, which are usually small due to their distance from the viewpoint. However, regions near the vanishing point may lack object, and not all objects may be sufficiently oversampled (Figure 4).
@@
Given an image, we aim to oversample object pixels, such as cars, and undersample background pixels, such as roads. Following [1, 35], we use kernel density estimation to model the bounding box saliency as a sum of Gaussians:
@@
where $c_{i}$, $w_{i}$ and $h_{i}$ represent the center, width, and height of the bounding box, $\mathcal{N}(\mu,\Sigma)$ is the normal distribution and $s$ is the saliency scale or bandwidth.
If many objects are small, we need to oversample objects more aggressively to mitigate background dominance. Conversely, if objects are already large enough, less aggressive oversampling is required. Therefore, we propose an expansion factor $f$ to determine object oversampling intensity:
@@
where $\Psi$ is the fraction of object size corresponding to an interval between size thresholds $t_{i}$ and $t_{i+1}$ (See Figure 1). In our case, we consider $\Psi=\{\psi_{small},\psi_{medium}$, $\psi_{large}\}$ to represent the percentage of small, medium, and large objects in the dataset, as per COCO [20] object size thresholds.
Saliency scale $s$ and expansion factor $f$ are inversely proportional ($s=\frac{P}{f}$), where $P$ is a constant. We clip $S$ between a lower bound of 0, to prevent negative saliency values that lead to holes, and an upper bound $U$, to discourage extreme warping.
When segmentation masks are available but bounding boxes are not, we convert the masks into ‘from-seg’ boxes. This does not significantly impact performance (See Section 5.3 and Appendix G).
Figure 6: Domain Adaptive Semantic Segmentation (Cityscapes $\rightarrow$ ACDC). Our warping significantly improved recognition of background elements like the sky, sidewalks, and buildings, as well as foreground objects such as traffic lights and traffic signs, indicating enhanced scene understanding ability.

### v3

# Instance-Warp: Saliency Guided Image Warping for Unsupervised Domain Adaptation

@@ Abstract
Driving is challenging in conditions like night, rain, and snow. Lack of good labeled datasets has hampered progress in scene understanding under such conditions. Unsupervised Domain Adaptation (UDA) using large labeled clear-day datasets is a promising research direction in such cases. However, many UDA methods are trained with dominant scene backgrounds (e.g., roads, sky, sidewalks) that appear dramatically different across domains. As a result, they struggle to learn effective features of smaller and often sparse foreground objects (e.g., people, vehicles, signs).
In this work, we improve UDA training by applying in-place image warping to focus on salient objects. We design instance-level saliency guidance to adaptively oversample object regions and undersample background areas, which reduces adverse effects from background context and enhances backbone feature learning. Our approach improves adaptation across geographies, lighting, and weather conditions, and is agnostic to the task (segmentation, detection), domain adaptation algorithm, saliency guidance, and underlying model architecture. Result highlights include +6.1 mAP50 for BDD100K Clear $\rightarrow$ DENSE Foggy, +3.7 mAP50 for BDD100K Day $\rightarrow$ Night, +3.0 mAP50 for BDD100K Clear $\rightarrow$ Rainy, and +6.3 mIoU for Cityscapes $\rightarrow$ ACDC. Besides, Our method adds minimal training memory and no additional inference latency. Code is available at https://github.com/ShenZheng2000/Instance-Warp.
@@ 1 Introduction
Figure 1: Object-Background Pixel Imbalance & Differences in Cross-Domain Object-Background Variations. Consider scenes from different domains: clear day, rainy, and night. red highlights foreground objects and yellow highlights background areas. Note, (a) background pixels occupy more space than foreground object pixels, (b) background elements like road exhibit higher cross-domain variations, while foreground objects like cars show smaller variations. Thus, focusing on objects would mitigate over-reliance on the background and improve domain adaptation.
@@
Camera-captured images, meanwhile, exhibit an object-background pixel imbalance, where pixel distribution is skewed towards backgrounds taking up more space. In contrast, objects are smaller (e.g. traffic signs), more scattered, occupying fewer pixels. The camera’s perspective further amplifies this effect, distant objects appear smaller and larger backgrounds to dominate the scene. However, many contemporary perception models treat image pixels, whether objects or backgrounds, uniformly. This leads to an over-reliance on the dominant backgrounds, making it challenging to learn effective features for less prevalent objects amid these dominant backgrounds.
@@
While this over-reliance on background pixels is a significant issue in traditional vision tasks, in Unsupervised Domain Adaptation (UDA) [15, 16, 19], which is crucial for driving in challenging conditions due to the scarcity of labeled data, the problem is further compounded by differences in cross-domain object-background variations, where background regions show greater variations than object regions across domains (Figure 2). Consequently, this over-reliance of UDA algorithms on background pixels makes adaptation difficult. Reducing dependence on these variable backgrounds would improve adaptation to new domains.
Figure 2: (a) Differences in Cross-domain object background variations. We visualize t-SNE [43] plots for ResNet-50 features extracted from ACDC [33] foggy and rainy images. Top plot shows image features (the majority of an image is background), while the bottom plot shows car features (the most common object). While image features across domains exhibit higher variance, car features are tightly entangled. Thus, over-reliance on the more variable background context makes adaptation difficult. (b) Saliency Guided Image Warping increases the relative size of salient foreground regions. Zoom in to see the car getting enlarged in the highlighted region, which reduces the effect of background context. (c) Learned Backbone features are better - they show a better focus on the object and less reliance on background.
@@
One possible remedy is data augmentation (e.g., resizing viz multi-scale training [36, 16] and inference [13, 35, 16]). However, these strategies introduce significant computational and storage overhead. Another approach, employing architectural priors to address these issues (e.g., feature pyramids [22, 25, 24, 20], coarse-to-fine feature alignment [54] or test-time scale optimization [9, 5]), only partially improve recognition with additional inference latency. Yet another strategy is to oversample or ‘zoom’ into salient image regions [27, 38, 8, 10, 39]. However, these methods do not distinguish between object and background, are only applicable to supervised settings, and introduce additional inference latency due to test-time operations.
@@
To address these limitations, our novel strategy judiciously utilizes available pixel space in an image to focus on salient regions for unsupervised domain adaptation. We adaptively oversample salient image regions corresponding to objects in-place, directly warping the original image without creating new images or additional crops. After that, we unwarp improved features while predicting labels and adapt models from a source to a target domain. This reduces background context dependence (Figure 2), significantly improving domain adaptability. During inference, we do not perform any warping as the learned features are already improved, thus adding no latency. We show significant improvements on real to real adaptation scenarios [15, 19, 16, 17] (Sections 4.14.2, Supp. A, Supp. B), including adaptation scenarios not tackled by prior works. In summary, our contributions are,
@@
  • We are the first to demonstrate that saliency guided image warping enhances backbone features and improves model adaptability for unsupervised domain adaptation (Section 4.2 and Supp. E).
@@
  • Our approach is efficient, with train-time image warping and feature unwarping but no test-time warping, resulting in minimal training overhead and no additional inference latency (Supp. B).
@@ 2 Related Work
Figure 3: Saliency Guided Image Warping for Unsupervised Domain Adaptation. Consider the standard UDA framework (See Sections 3.1 and 3.4) with supervised pre-training and unsupervised self-training phases (as shown here). The components of warping and unwarping are marked in cyan. We warp images using saliency guidance to oversample salient image regions, encouraging improved feature learning for the backhone. Our instance-level saliency guidance oversamples object regions, showing better performance compared to Static Prior [38, 39] and Geometric Prior [10]. We unwarp features before predicting labels, ensuring that our labels are never warped, and the UDA losses of the employed algorithm (e.g., [15, 19, 16, 17, 21, 7, 53]) remain unmodified. We do not warp or unwarp at test time.
@@
Saliency Guided Image Warping. Non-uniform Spatial Transformations were traditionally used to correct image distortions arising from camera acquisition [47, 1]. Recently, spatial transformation methods [18, 27] for learned models introduce intentional image distortions to enhance performance. Learning to Zoom [27] that over-sampling task-specific salient regions is useful, as visual recognition requires combining information from different spatial resolutions. Building on this paradigm, warping for other tasks like detection [38, 39] and segmentation [39] has been explored. For e.g., Fovea [38] and Two Plane Prior [10] designed ‘salient’ priors to guide warping, introducing static & temporal, and geometric saliency guidance, respectively.
@@
Our approach stands outs in the following ways: (a) Unlike many UDA methods that indiscriminately process object and background, we introduce an instance-level saliency guidance that targets regions containing objects. (b) Unlike previous image warping efforts focused on supervised settings, our novel training strategy and saliency guidance tailor image warping for unsupervised domain adaptation. (c) Unlike prior UDA works using augmentations such as resizing [16], which are memory-intensive, or supervised warping works [38, 10, 39] that perform test-time warping and increase inference latency, our in-place warping incurs minimal memory and latency overhead during training and no additional latency during testing.
@@
Image Transformations for Domain Adaptation. Transformations like flipping, cropping, blurring, color jittering, grayscaling, and contrast adjustment are common perturbations in domain adaptation [41, 50, 51, 26, 32, 53, 40, 7, 45, 48, 21, 15, 16, 14, 17, 19]. They increase source domain diversity and narrow the gap to target domain. For example, 2PCNet [19] introduces a domain-specific Night Augmentation (NightAug) to make day images resemble night images. To the best of our knowledge, in-place image warping as a domain adaptive perturbation has not been explored previously. Our approach is agnostic to domain adaptation algorithms and can be seamlessly combined with other augmentation strategies.
@@ 3 Method
3.1 Overview
We perform unsupervised domain adaptation (UDA), i.e. source domain $\mathcal{X}_{a}$ includes labeled data, i.e., $X_{a}=\{\ (x_{a},y_{a})\in\mathcal{X}_{a}\ \}$ and target domain $\mathcal{X}_{b}$ includes unlabeled images, i.e., $X_{b}=\{\ x_{b}\in\mathcal{X}_{b}\ \}$.
The core of our approach lies in the operations of “image warping” and “feature unwarping”. In a standard UDA framework [15, 19], self-training involves a student and teacher network trained on different image augmentations or views [4, 11] with an unsupervised loss following a supervised pre-training phase. In our method, the model views “warped” images to encourage feature learning in oversampled regions. Although our model processes warped images, it predicts labels as if the image was never warped, accomplished through feature “unwarping” [39]. This ensures that ground truth or teacher labels remain unwarped, allowing unsupervised losses (e.g., from [15, 19, 16, 17, 21, 7, 53]) to be used without modification. The “warp” and “unwarp” operations are employed during domain adaptation to improve backbone feature learning. The operations are not applied at test time, ensuing no added inference latency. Figure 1 illustrates the mechanism of Saliency Guided Warping for domain adaptation.
In Section 3.2, we discuss how to warp and unwarp images. Section 3.3, we discuss strategies of deciding which regions to sample to encourage improved feature learning. Section 3.4 discusses how to employ the pair of operations during domain adaptation.
@@ 3.2 Image Warping and Feature Unwarping
Image Warping. Following common practice in saliency guided image warping [1, 27, 38] where warp $\mathcal{T}$ is defined through a backward map, we consider inverse transformation $\mathcal{T}^{-1}_{S}$ parameterized by saliency $S$ such that,
@@
a resampled $I^{\prime}$ for every location $\mathbf{u}$ is obtained by using warp $W_{\mathcal{T}}$ given input image $I$ and saliency guidance $S$, which is a 2D map of the output image size. The rationale and choice of $S$ is described in Section 3.3. Warping incurs an additional 1.4 ms latency per image on a 4090 GTX GPU and has no additional learned parameters.
Feature Unwarping. If we operate in the space of warped images, our predictions would also be warped. However, we do not wish to warp the labels to match the warped input image. Instead, we perform feature unwarping on the spatial backbone features and train the model without warping labels or unwarping predictions like [38, 10], and the losses remain unchanged. This makes our warping task-agnostic. However, $\mathcal{T}^{-1}$ has no closed form inverse [38, 39]. Therefore, we adopt the strategy from LZU [39], which approximates $\mathcal{T}^{-1}$ with $\widetilde{\mathcal{T}}^{-1}$, a piecewise tiling of invertible bilinear maps, ensuring that both $\widetilde{\mathcal{T}}$ and $\widetilde{\mathcal{T}}^{-1}$ exist):
${\mathcal{T}}(\mathbf{u})\approx\widetilde{\mathcal{T}}(\mathbf{u})=\begin{cases}\widetilde{\mathcal{T}}_{ij}(\mathbf{u})&\text{ if }\mathbf{u}\in\operatorname{Range}\left(\widetilde{\mathcal{T}}_{ij}^{-1}\right)\\ 0&\text{ else }\end{cases}$ (2)
We then unwarp features using the approximation $\widetilde{\mathcal{T}}$ (the closed-form inverse of $\widetilde{\mathcal{T}}^{-1}$) as our backward map, since the roles of output and input coordinates are switched during unwarping. Unwarping incurs 4.2 ms latency per image on a 4090 GTX GPU and does not introduce any additional learned parameters. Figure 5 shows that while warping and unwarping are lossy, the information loss is minimal.
Additional Considerations. We restrict the space of allowable warps for $\mathcal{T}^{-1}$ to preserve axis alignment. In this approach, entire rows or columns are ‘stretched’ or ‘compressed,’ which are shown by [38, 10, 39] to perform better,
[…]
where $k_{u}$ and $k_{v}$ are Gaussian kernels and $\mathbf{u}=(u,v)$ represent the two image axes in pixel space. $S_{u}$ and $S_{v}$ are the marginalized versions of $S$ along $u$ and $v$ axes, respectively.
@@ 3.3 Designing Saliency Guidance
Figure 4: Image Warping with Different Saliency Guidance Functions. In-place warping follows a zero-sum pixel constraint: enlarging one region necessitates shrinking another. Bounding boxes mark small, medium, and large objects. Our instance-level saliency guidance oversamples objects and undersamples the background. In contrast, Static Prior fails when object locations do not align with the dataset’s average object location, while Geometric Prior fails for small objects not near the vanishing point.
Saliency ($S$) determines which regions of an image are sampled more during warping [27]. We explore three saliency functions: two based on approximate guidance from prior work [38, 39, 10], and a novel instance-level saliency guidance that we propose. Since images are warped in-place, expanding some regions necessitates reducing others (Figure 4). Only instance-level saliency guidance oversamples objects and undersamples the background.
Approximate Saliency Guidance. The Geometric Prior [10] saliency guidance is a function of the dominant vanishing point in the scene, oversampling objects close to the vanishing point, which are usually small due to their distance from the viewpoint. However, regions near the vanishing point may lack object, and not all objects may be sufficiently oversampled (Figure 4).
@@
Given an image, we aim to oversample object pixels, such as cars, and undersample background pixels, such as roads. Following [2, 38], we use kernel density estimation to model the bounding box saliency as a sum of Gaussians,
Figure 5: Warping and Unwarping. Shown is a warped image with two different saliency scales $s=\{1,16\}$. A higher saliency scale implies less intense warping. While unwarping is applied to features, here we apply it to the warped image for illusration. Although saliency guided warping and unwarping are lossy, we observe that the error between the original image and warped-then-unwarped image is very low, indicating minimal information loss.
@@
where $c_{i}$, $w_{i}$ and $h_{i}$ represent the center, width, and height of the bounding box, $\mathcal{N}(\mu,\Sigma)$ is the normal distribution and $s$ is the saliency scale. Figure 5 shows the effect of different saliency scales, where a higher scale results in more warping areas but less warping intensity. The expansion factor $f$ ($s=\frac{P}{f}$) reflects warping intensity, where $P$ is a constant.
Intuitively, for dataset with many small objects, high warping intensity is helpful. However, if the dataset contains mostly large objects, warping would distort them and harm performance. In this case, the warping intensity should be reduced. We define $f$ following this intuition as:
@@
where $\Psi$ is the fraction of object size corresponding to an interval between size thresholds $t_{i}$ and $t_{i+1}$. In our case, we consider $\Psi=\{\psi_{small},\psi_{medium}$, $\psi_{large}\}$ to represent the percentage of small, medium, and large objects in the dataset, as per COCO [23] object size thresholds. In the rare case of having only large objects in the dataset, $f$ approaches 1 (minimal warping). Conversely, with only small objects, $f$ tends to a large value (very intense warping).
We clip $S$ between a lower bound of 0, to prevent negative saliency values that lead to holes, and an upper bound $U$, to discourage extreme warping. Figure 5 shows that a higher saliency scale $s$ results in less intense warping.
When segmentation masks are available but bounding boxes are not, we convert the masks into ‘from-seg’ boxes. This does not significantly impact performance (See Section 4.3 and Supp. G).
Figure 6: Domain Adaptive Object Detection (BDD100K Day $\rightarrow$ Night). Our warping improves scene understanding by enabling 2PCNet to better distinguish between different objects and between objects and backgrounds. In contrast, 2PCNet [19] baseline misclassifies roads as cars and streetlights as traffic signs.
Figure 7: Domain Adaptive Semantic Segmentation (Cityscapes $\rightarrow$ ACDC). Our warping improved recognition of backgrounds (e,g., sky, sidewalks, and buildings), and foreground objects (e.g., traffic lights, traffic signs), indicating enhanced scene understanding.